a proton moving along the x axis moves with a constant accelaration of 6 x 10^12 m/s with initial velocity of 4x10^6 m/s . what is the velocity of proton after it travels a distance of 80cm?

Grade:12

1 Answers

Aman Bansal
12 years ago

Dear Goutham,

using linear equations of motion
v^2-u^2=2as
given a=6*1012m/s^2~i-i~s=0i-i8m~i-i~u=4*10^6m/s
v^2=2*0i-i8*6*10^12 +(4*10^6)^2
=(9i-i6+16)*10^12
=25i-i6*10^12
thus final velocity is sqrt(25i-i6*10^12)m/m

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert